Data from the Office for National Statistics show that the mean age at which men in Great Britain get married was 33.0. A news reporter noted that this represents a continuation of the trend of waiting until a later age to wed. A new sample of 47 recently wed British men provided their age at the time of marriage. These data are contained in the Excel Online file below. Construct a spreadsheet to answer the following questions. Open spreadsheet Do these data indicate that the mean age of British men at the time of marriage exceeds the mean age in 2013? Test this hypothesis at . What is your conclusion? Use the obtained rounded values in your calculations. Sample mean: 33.11 years (to 2 decimals) Sample standard deviation: 4.7468 years (to 4 decimals) -value: 0.154 (to 3 decimals) -value: (to 3 decimals) Because -value , we . There is evidence to conclude that the mean age at which British men get married exceeds what it was in 2013. Excel giving me wrong p value.

Find me the P- Value

Data set

25

30

30

33

34

34

38

37

29

39

30

29

37

26

30

34

39

26

28

34

35

40

25

35

37

31

35

25

40

40

32

34

35

28

38

37

39

31

37

37

39

34

39

30

25

29

27
Answer #1

Solution:

Here, we have to use one sample t test for population mean.

H0: µ = 33 versus Ha: µ > 33

This is an upper tailed or right tailed test.

From given data, we have

Xbar = 33.11

S = 4.7468

n = 47

df = n – 1 = 46

We assume α = 0.05

Test statistic formula is given as below:

t = (Xbar - µ)/[S/sqrt(n)]

t = (33.11 – 33)/[4.7468/sqrt(47)]

t = 0.11/ 0.6924

t = 0.1589

P-value = 0.4372

(by using t-table or excel)

P-value > α = 0.05

So, we do not reject the null hypothesis

There is insufficient evidence to conclude that the mean age at which British men get married exceeds what it was in 2013.